# IDP-Bert: Predicting Properties of Intrinsically Disordered Proteins (IDP) Using Large Language Models

### v1

@@ 4 Conclusion
This study introduces IDP-Bert model, a fine-tuned protein language model (PLM), to effectively predict Intrinsically Disordered Protein (IDPs) properties. Given the inherent lack of well-defined structures in IDPs, this model is designed to decipher the language of IDPs and precisely predict their intrinsic characteristics based solely on amino acid sequences. Through our experiments, accurate predictions of structural, dynamic, and thermodynamic properties, including Radius of Gyration, end-to-end Decorrelation Time, and Heat Capacity, were achieved. Additionally, analysis of attention weights assigned to individual amino acids within IDPs illuminates their contributions to determining protein properties that facilitate protein engineering and drug design efforts.
@@ 5 Data and software availability
The necessary information containing the codes and data for downstream tasks used in this study is available here: https://github.com/ChakradharG/IDP-BERT

### v2

@@ 4 Conclusion
This study introduces IDP-Bert model, a fine-tuned protein language model (PLM), to effectively predict some of the Intrinsically Disordered Protein (IDPs) properties. Given the inherent lack of well-defined structures in IDPs, this model is designed to decipher the language of IDPs and precisely predict their intrinsic characteristics based solely on amino acid sequences. Through our experiments, accurate predictions of structural, dynamic, and thermodynamic properties, including Radius of Gyration, end-to-end Decorrelation Time, and Heat Capacity, were achieved. Additionally, analysis of attention weights assigned to individual amino acids within IDPs illuminates their contributions to determining protein properties that facilitate protein engineering and drug design efforts.
@@ 5 Data and software availability
The necessary information containing the codes and data for downstream tasks used in this study is available here: https://github.com/DanushSadasivam/IDP-BERT
6 Supporting Information file
Experimental details, including six experiments with varying hyperparameters such as hidden layer size, number of attention heads, and the number of hidden layers, and their corresponding R² values for train, validation, and test sets. Analysis of the contribution of sequence length in predicting properties, with modified training procedures and the resulting performance metrics. Protein taxonomy data obtained from a similarity search on the UniProt database, including the UniProt IDs of the similarity hits, similarity fractions, and taxonomy information in a CSV file named ”taxonomy results.csv”.